OpenStudy (anonymous):

which one is bigger 19/42 or 13/14

OpenStudy (anonymous):

13/14

OpenStudy (diyadiya):

13/14

OpenStudy (anonymous):

19/42 is equivalent to 0.45238095... and 13/14 is equivalent to 0.92857142857 so 13/14 is greater or so 19/42<13/14

OpenStudy (jamesj):

The best way to answer this question is to subtract the two and see if the answer is negative or positive.\[\frac{19}{42} - \frac{13}{14} =\frac{19}{42} - \frac{13 \times 3}{14 \times 3} = \frac{19 - 39}{42} =\frac{-20}{42}\] Hence 19/42 must be less than 13/14 because \[\frac{19}{42} - \frac{13}{14} < 0\]